### v1

@@ Abstract
Figure 1: From a set of monocular images capturing a moving scene, a dynamic 3D representation is learned using time-varying Gaussians (a). Then the articulated parts (depicted with the trajectories of the motion) are identified using masking (b). This allows for learning a fine-scale, per-Gaussian level of control (c). The approach is capable of synthesizing novel configurations not present in the original sequence, for example, independently opening the hood and trunk of the toy car.
@@ 2.3 Gaussian Splatting
Figure 2: CoGS Overview. CoGS consists of two parts: Dynamic GS and Controllable GS.
@@ 3.1 Dynamic Gaussian Splatting
To represent dynamic scenes and ultimately introduce fine-scale attribute control with 3D Gaussians, we utilize the differentiable Gaussian rasterization pipeline proposed by [14]. In specific, we directly follow the method therein and then augment its static properties with deformation fields to model scene dynamics.
@@ 3.1.2 Optimization for Dynamic Scene Representation
where $t$ is the current time step. Different from [42], we also learn an offset for color to account for any changes that may occur over time (e.g. shadows and reflections). The output from these networks is then added to the corresponding parameters, and the differentiable rasterization pipeline proceeds.
@@
Demonstrated in Fig. 3(a), this loss yields a much more consistent dynamic representation when compared to without it as shown in 3(b).
@@ 3.2 Controllable Gaussian Splatting
Addressing these challenges, we propose an effective method to obtain the mask vector $m$. We start by acquiring $K$ 2D masks for the 2D frames, where $K$ can vary from all frames (for scenarios with available automatic mask generation methods like face recognition [45], as illustrated in Fig. 6(d)) to a single frame (which can be manually labeled, as demonstrated in Figs. 6(a) & 6(b) & 6(c)). After the 2D mask acquisition, we perform a 2D-to-3D mask projection. A practical method involves associating the 3D point with the corresponding 2D pixel, utilizing depth maps and camera poses as in [22]. However, this method falls short of attaining the fine-grained boundary of the controllable part due to the splatting process 3.1.1.
@@ 4.1 Datasets
Toy Car Scene. We also ran experiments on a self-captured Toy Car scene. The capture process involved manually opening the car’s trunk and hood separately, recording one video per transition. Once the transitions were completed, we stitched the individual videos together to create a single cohesive dynamic scene. This scene was captured with an Apple iPhone 13 Pro.
@@ 4.3 Results
Table 3: Quantitative results on controllable face scenes. We color code each row as best, second best, and third best.
[…]
NeRF [24] 28.80 0.951 21.0
NeRF + Latent [24] 32.65 0.981 18.2
Nerfies[29] 32.27 0.981 18.0
HyperNeRF[30] 32.52 0.981 16.9
CoNeRF-$M$[13] 32.06 0.979 16.7
CoNeRF[13] 32.34 0.981 16.8
3D GS [14] 25.94 0.834 41.4
Ours 32.60 0.983 16.4

### v2

@@ Abstract
Figure 1: From a set of monocular images capturing a moving scene, a dynamic 3D representation is learned using time-varying Gaussians (a). Then the articulated parts (depicted with the trajectories of the motion) are identified using masking (b). This allows for learning a fine-scale, per-Gaussian level of control (c). The approach is capable of synthesizing novel configurations not present in the original sequence, for example, independently opening the hood, trunk, and doors of the toy car.
@@ 2.3 Gaussian Splatting
Figure 2: CoGS Overview. CoGS consists of two parts: Dynamic GS and Controllable GS. For Dynamic GS, an offset is learned for $(\mu,C,R,S)$ by separate MLPs (only one shown in figure). To extend to controllable scenarios, signals extracted from the dynamic model are used to obtain attribute offsets, which are then masked to affect the desired control region.
@@ 3.1 Dynamic Gaussian Splatting
To represent dynamic scenes and ultimately introduce fine-scale attribute control with 3D Gaussians, we utilize the differentiable Gaussian rasterization pipeline proposed by [14]. Specifically, we directly follow the method therein and then augment its static properties with deformation fields to model scene dynamics.
@@ 3.1.2 Optimization for Dynamic Scene Representation
where $t$ is the current time step. Different from [42], we also learn an offset for color to account for any changes that may occur over time (e.g. shadows and reflections). The outputs from these networks are then added to the corresponding parameters, and the differentiable rasterization pipeline proceeds.
@@
Demonstrated in Fig. 4(a), this loss yields a much more consistent dynamic representation when compared to without it as shown in 4(b).
@@ 3.2 Controllable Gaussian Splatting
Addressing these challenges, we propose an effective method to obtain the mask vector $m$. We start by acquiring $K$ 2D masks for the 2D frames, where $K$ can vary from all frames (for scenarios with available automatic mask generation methods like face recognition [45], as illustrated in Fig. 7(d)) to a single frame (which can be manually labeled, as demonstrated in Figs. 7(a) & 7(b) & 7(c)). After the 2D mask acquisition, we perform a 2D-to-3D mask projection. A practical method involves associating the 3D point with the corresponding 2D pixel, utilizing depth maps and camera poses as in [22]. However, this method falls short of attaining the fine-grained boundary of the controllable part due to the splatting process 3.1.1.
@@ 4.1 Datasets
Toy Car Scene. We also ran experiments on two self-captured Toy Car scenes. The capture process involved manually opening the regions of control (doors, hood, trunk, etc.) and recording one video per transition. Once the transitions were completed, we stitched the individual videos together to create a single cohesive dynamic scene. This scene was captured with an Apple iPhone 13 Pro.
@@ 5 Conclusion
Our method is not without limitations. CoGS faces challenges with shiny or intricately lit objects, common in GS pipelines. Dynamic modeling may struggle with non-rigid deformation and large-scale movements in monocular settings. Limitations may also arise from controllable signal extraction and re-alignment, with the current PCA method potentially struggling with highly complex movements. Addressing these limitations will be the focus of future work.
Acknowledgements
This research was supported partially by Fujitsu.
Supplementary Material
6 Overview
This supplementary material offers comprehensive quantitative data and further qualitative insights, highlighting the advantages of our newly developed Dynamic 3D Gaussian Splatting (GS) and Controllable GS methods. In addition, we have included illustrative videos on the attached webpage, providing a dynamic visual representation of our methods in action.
7 Per-Scene Quantitative Results
For completeness, we provide detailed per-scene quantitative results for reconstruction quality metrics, including PSNR, SSIM, and LPIPS. These are presented for both synthetic (Tab. 3) and real (Tab. 4) dynamic scenes. This extension to Tables 1 and 2 from the main paper offers a more nuanced view, as it disaggregates the average performance metrics across different scenes. Our analysis reveals that our Dynamic GS method exhibits superior performance in synthetic scene datasets while achieving comparable results in real scenes. This difference in performance might be attributed to the challenges inherent in modeling the movement of Gaussians using a single camera setup.
8 More Qualitative Results
Additional qualitative results can be found on the project’s website (https://cogs2024.github.io). For optimal viewing, please open the link using the Chrome browser.
Table 3: Per-scene quantitative results on synthetic dynamic scenes. We color code each row as best, second best, and third best.
Hell Warrior Mutant Hook Bouncing Balls
[…]
NeRF[24] 13.52 0.8100 0.2500 20.31 0.9100 0.0900 16.65 0.8400 0.1900 20.26 0.9100 0.2000
DirectVoxGo[34] 13.51 0.7500 0.2500 19.45 0.8900 0.1200 16.16 0.8000 0.2100 20.20 0.8700 0.2200
Plenoxels[8] 15.19 0.7800 0.2700 21.44 0.9100 0.0900 17.90 0.8100 0.2100 21.30 0.8900 0.1800
T-NeRF[32] 23.19 0.9300 0.0800 30.56 0.9600 0.0400 27.21 0.9400 0.0600 37.81 0.9800 0.1200
D-NeRF[32] 25.10 0.9500 0.0600 31.29 0.9700 0.0200 29.25 0.9600 0.1100 38.93 0.9800 0.1000
TiNeuVox-S[7] 27.00 0.9500 0.0900 31.09 0.9600 0.0500 29.30 0.9500 0.0700 39.05 0.9900 0.0600
TiNeuVox-B[7] 28.17 0.9700 0.0700 33.61 0.9800 0.0300 31.45 0.9700 0.0500 40.73 0.9900 0.0400
3D GS [14] 29.72 0.9129 0.1215 23.59 0.9318 0.0631 21.88 0.8847 0.1104 23.03 0.9583 0.0737
Ours 40.43 0.9812 0.0267 42.14 0.9937 0.0063 36.43 0.9838 0.0174 40.98 0.9958 0.0103
Lego T-Rex Stand Up Jumping Jacks
[…]
NeRF [24] 20.30 0.7900 0.2300 24.29 0.9300 0.1300 18.19 0.8900 0.1400 18.28 0.8800 0.2300
DirectVoxGo [34] 21.13 0.9000 0.1000 23.27 0.9200 0.0900 17.58 0.8600 0.1600 17.80 0.8400 0.2000
Plenoxels [8] 21.97 0.9000 0.1100 25.18 0.9300 0.0800 18.76 0.8700 0.1500 20.18 0.8600 0.1900
T-NeRF [32] 23.82 0.9000 0.1500 30.19 0.9600 0.1300 31.24 0.9700 0.0200 32.01 0.9700 0.0300
D-NeRF [32] 21.64 0.8300 0.1600 31.75 0.9700 0.0300 32.79 0.9800 0.0200 32.80 0.9800 0.0300
TiNeuVox-S [7] 24.35 0.8800 0.1300 29.95 0.9600 0.0600 32.89 0.9800 0.0300 32.33 0.9700 0.0400
TiNeuVox-B [7] 25.02 0.9200 0.0700 32.70 0.9800 0.0300 35.43 0.9900 0.0200 34.23 0.9800 0.0300
3D GS [14] 22.73 0.9282 0.0679 21.92 0.9537 0.0498 21.54 0.9283 0.0854 20.16 0.9279 0.0855
Ours 25.16 0.9451 0.0421 37.25 0.9923 0.0115 43.35 0.9929 0.0092 37.48 0.9891 0.0158
Table 4: Per-scene quantitative results on real dynamic scenes. We color code each row as best, second best, and third best.
torchocolate cut-lemon chickchicken hand
[…]
NeRF [24] 22.5 0.866 0.373 24.1 0.826 0.437 18.8 0.761 0.453 23.8 0.773 0.469
NV [21] 24.6 0.917 0.189 28.8 0.951 0.190 22.6 0.861 0.243 29.3 0.912 0.213
NSFF [19] 22.3 0.883 0.253 28.0 0.904 0.238 27.7 0.939 0.173 24.9 0.797 0.329
Nerfies [29] 27.8 0.959 0.169 30.8 0.946 0.223 28.7 0.948 0.141 29.9 0.940 0.171
HyperNeRF [30] 28.0 0.962 0.172 31.8 0.956 0.210 28.7 0.948 0.156 30.7 0.950 0.150
TiNeuVox-S [7] 21.5 0.754 0.478 23.4 0.642 0.604 25.3 0.761 0.485 24.2 0.604 0.614
TiNeuVox-B [7] 27.1 0.824 0.395 28.6 0.694 0.509 29.0 0.812 0.408 27.3 0.678 0.493
3D GS [14] 21.8 0.787 0.402 22.6 0.667 0.482 20.2 0.719 0.517 23.6 0.723 0.351
Ours 28.3 0.949 0.174 31.4 0.945 0.205 28.8 0.942 0.146 30.8 0.947 0.161